OpenStudy (anonymous):

Which of the following are solutions to the quadratic equation? 2x2 + 7x - 2 = - x2 + 2x A. 2 B. 1 C. -2 D. 3 E. F. -

OpenStudy (gorv):

2x2+x2+7x-2x-2=0 3x2+5x-2=0 3x2+6x-x-2=0 3x(x+2)-1(x+2)=0 (x+2)(3x+2)=0 x=-2 0r -2/3